13.Electron are accelerated in a television tube through a potential difference of 9.8kV. Find the highest frequency and minimum wavelength of the electromagnetic wave emitted, when these strike the screen of the tube. In which region of the spectrum will these lie? a. Differentiate X-rays and visible light in terms of the origin of these EM radiations. b. What voltage must be applied to an X-ray tube for it to emit X-rays with minimum wavelength 0.5 A˚ ?


1
Expert's answer
2021-12-10T11:31:07-0500

"E=A,"

a)

"h\nu=eU,\implies"

"\nu=\frac{eU}h=3.37\cdot10^{18}~Hz,"

"\lambda=\frac c{\nu}=0.89~Å,"

b)

"\frac{hc}{\lambda}=eU',\implies"

"U'=\frac{hc}{e\lambda}=24.86~kV."